1. 60 ml. =    oz
2. 8oz. = ml
3. 2tsp= cc
4. labeled 125mg per 5ml. The order is for 0.25Gm Give in ml.
5. Give Sumycin 4Gm stat. It's available in 500mg tablets. Give in tablets.
6. 7.5 mg per kg. your client's weight is 88kg. Give in mg
7. Heparin 8,000 units subcutaneously TID. The heparin comes in a solution of 2,500 units/ml. How many milliliter should the nurse administer to the patient?
8. 5mg per cc. How much is needed to give 5mg dose? ml
Answer #1

1. In the US, 1 ounce is given by

1 fluid ounce = 29.574 ml.

Therefore, 1 ml = 1/29.574 oz

Thus, 60 ml = 60 x 1/29.574 oz

= 2.029 oz

Hence, 60 ml = 2.029 oz.

2. Again,

1 fluid ounce = 29.574 ml.

Therefore, 8 oz = 8 x 29.574 ml

= 236.592 ml

Hence, 8 oz = 236.592 ml.

3. In the US, 1 teaspoon contains 4.929 cubic centimeter volume of liquid.

Therefore, 2 tsp = 2 x 4.929 cc

= 9.858 cc.

Hence, 2 tsp = 9.858 cc.

4. Given drug is in the concentration = 125 mg/5 ml.

Order is for = 0.25 gm

Now, 1 gm = 1000 mg.

Therefore, 0.25 gm = 0.25 x 1000 mg

= 250 mg.

Now, according to the question, 125 mg drug is present in 5 ml.

Therefore, 1 mg drug is present in 5/125 ml.

Now, according to the order, 250 mg drug is present in 250 x 5/125 ml = 10 ml.

Therefore, 0.25 gm is present in 10 ml.

5. Available units of Sumycin = 500 mg tablets.

Given order of Sumycin = 4 gm

= 4 x 1000 mg

= 4000 mg.

Now, 1 tablet has 500 mg of Sumycin.

Therefore, to administer 4000 mg of Sumycin, no. of tablets required = 4000 mg/500 mg = 8

Therefore 8 tablets are required to administer 4 gm of Sumycin.

6. Required dosage = 7.5 mg/kg.

Weight of client = 88 kg.

Therefore, total dosage required = 88 kg x 7.5 mg/kg.

= 660 mg.

Hence, 660 mg dosage is required.

7. According to the question, Heparin should be administered subcutaneously 8000 units thrice in a day(TID).

Available solution of heparin = 2500 units/ml.

Thus, 1 ml contains 2500 units of Heparin.

Therefore, volume required for 8000 units of Heparin = 8000/2500 = 3.2 ml.

Hence, the nurse should administer 3.2 ml subcutaneously thrice a day to deliver the required dose to the patient.

8. According to the question, 5 mg is present in 1 cc.

As is known, 1 cubic centimeter = 1 ml

Therefore, 5 mg is present in 1 cc = 1 ml.

Hence, to give a 5 mg dose, 1 ml needs to be administered.